It is known that pancake sales at an athletics meeting between a number of schools follow the following pattern: when the price is R3,00 per pancake, 3 000 pancakes are sold during the day and when the price is R2,50 each, 5 000 are sold. Find the demand function with quantity, Q, as a function of price, P. [1] Q = −4P + 15 [2] Q = −4 000P + 15 000 [3] Q = −1 000P + 6 000 [4] Q = 4 000P − 9 000 


1
Expert's answer
2021-06-01T08:08:52-0400

Demand function:

"Q=a-bP"


We have:

For Q=3000, P=3.00:

"3000=a-3b"

For Q=5000, P=2.50:

"5000=a-2.5b"


Then:

"2000=0.5b"

"b=4000"

"a=3000+3\cdot4000=15000"


Answer: [2] Q = −4 000P + 15 000